Find out the financial position of your family by calculating the Income and Expenditure

data. Clearly mention if you find any savings? Discuss the possibilities of Loan in case of

excess expenditures?


1
Expert's answer
2021-04-20T12:33:45-0400

Dept of public health ($) Family business($) Total

Salaries 15000 50000 65000

Fringe 200 3000 4000

Food stuff 1500 2000 3500

phone 500 1000 1500

wifi 300 500 800

insurance 400 800 1200

rent 1500 4000 5500

Total 6700 61300 81500


Expenditure: 4000+3500+1500+800+1200+5500= $16500

Income : $65000

        65000-16500= $48500

 Savings: 4000

     48500+4000= $52500

There is no need for loan